0 Required information The following information applies to the questions displayed below. Prat Corp. started the 2018 accounting period with $32,000 of assets (all cash). $13,000 of liabilities, and $7,000 of common stock. During the year, the Retained Earnings account increased by $9,550. The bookkeeper reported that Prat paid cash expenses of $27,000 and paid a $2,200 cash dividend to the stockholders, but she could not find a record of the amount of cash that Prat received for performing services. Prat also paid $7,000 cash to reduce the liability owed to the bank, and the business acquired $7,200 of additional cash from the issue of common stock Required (Hint Determine the amount of beginning retained earnings before considering the effects of the current period events. It also might help to record all events under an accounting equation before preparing the statements.) a-1. Prepare an income statement for the 2018 accounting period. a-2. Prepare a statement of changes in stockholders equity for the 2018 accounting period. a-3. Prepare a period-end balance sheet for the 2018 accounting period. a-4. Prepare a statement of cash flows for the 2018 accounting period. Complete this question by entering your answers in the tabs below. Req A1 Req A2 Req A3 Req A4 Prepare an income statement for the 2018 accounting period Prat Corp. Prev 1 2 of 9 l Next >
Answer #1
Beginnig retained earnigs:
Assets=Liabilities+ Stockholder's equity
32000=13000+Stockholder's equity
Stockholder's equity=32000-13000=$ 19000
Stockholder's equity=common stock+retained earnings
19000=7000+retained earings
Retained earnings=19000-7000=$ 12000
Cash received from performing services:
Let's assume that increase in retained earnigs=Net income=$ 9550
Cash expenses=$ 27000
Net income=Sales-expenses
9550=Sales-27000
Sales=9550+27000=$ 36550
a-1. Income statement
$
Sales 36550
Less: Expenses 27000
Net income 9550
a-2. Statemet of chages in stockholder's equiry
Common
stock
Retained
earnings
Beginning balance 7000 12000
Net income 9550
Dividend -2200
Issue of common stock 7200
Ending balace 14200 19350
a-3. Balance sheet
$
Assets:
Cash (Note:1) 39550
Total assets 39550
Liabilities and stockholder's equity
Liabilities (13000-7000) 6000
Stockholder's equity:
Common stock 14200
Retained earnigs 19350
Total stockholder's equity 33550
Total liabilities and stockholder's equity 39550
Note:1
Cash
$ $
Beginning balance 32000
Add:
Cash sales 36550
Issue of common stock 7200 43750
75750
Less:
Expenses 27000
Dividend 2200
Liability reduced 7000 36200
Ending balance 39550
a-4. Statemet of cashflows
$ $
Cashflow from operating activities:
Net income 9550
Adj:
Non-cash expenses 0
Change in current assets and liabilities 0 0
Net cash from operating activities 9550
Cashflow from investing activities 0
Cashflow from financing activities:
Dividend paid -2200
Cash paid to reduce liability -7000
Issue of common stock 7200
Net cash used in financing activities -2000
Net increase in cash 7550
Add: Beginning balance of cash 32000
Ending balance of cash 39550